OpenStudy (anonymous):

Find the absolute value of each complex number. |7 - i|

OpenStudy (anonymous):

I know the answer is 5sqrt(2), but I don't know how they got that

jimthompson5910 (jim_thompson5910):

|a+bi| = sqrt(a^2+b^2) |7+(-1)i| = sqrt(7^2+(-1)^2) I'll let you take it from here

OpenStudy (anonymous):

why did you get rid of the i in sqrt(7^2+(-1)^2)?

jimthompson5910 (jim_thompson5910):

You use the formula |a+bi| = sqrt(a^2+b^2) In the case of 7 - i, it's in the form 7+(-1)i which means a = 7 and b = -1

jimthompson5910 (jim_thompson5910):

So you plug those values into the formula |a+bi| = sqrt(a^2+b^2)

OpenStudy (anonymous):

Wait I don't understand... isn't i = sqrt(-1)?

jimthompson5910 (jim_thompson5910):

yes, but with that formula, you only worry about the coefficients

OpenStudy (anonymous):

What do you do with the i?

jimthompson5910 (jim_thompson5910):

If you plot 7 - i in the complex number plane, you're basically plotting the point (7,-1) on the xy coordinate system To find |7-i|, you're finding the distance from (0,0) to (7,-1)

jimthompson5910 (jim_thompson5910):

So you can see that finding the distance from (0,0) to (7,-1) doesn't involve the term i

OpenStudy (anonymous):

ooh

jimthompson5910 (jim_thompson5910):

glad things are clicking

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

yw